OpenStudy (yoloshroom):

Estabon poured himself a hot beverage that had a temperature of 205F then set it on the kitchen table to cool. The temp of the kitchen ws a constant 71F. If the drink cooled to 182 F in 3 min, how long will it take to cool to 105 F???

OpenStudy (yoloshroom):

@AlexandervonHumboldt2

OpenStudy (yoloshroom):

@Luigi0210

OpenStudy (triciaal):

hint y = m x + b y is the hot temp m is the drop in temp per min and b is the temp of the room

OpenStudy (yoloshroom):

Doesnt this problem pertain to newton's law of cooling?

rvc (rvc):

rate of cooling? do you know the formula?

OpenStudy (baru):

hi :D

OpenStudy (yoloshroom):

nu ;-;

rvc (rvc):

d theta/dt

OpenStudy (baru):

newtons law of cooling: rate of change of temprature of a body is proportional to difference between its temprature and surrounding temperature

rvc (rvc):

perfect :)

OpenStudy (baru):

\(T_b\): temperature of beverage \(T_k\): temperature of kitchen=71 \[\frac{dT_b}{dt}=-k(T_b-T_k)\]

OpenStudy (baru):

substitute \(T_k\) ; solve the differential equation for \(T_b\)

OpenStudy (yoloshroom):

annnd im still confused lol

OpenStudy (baru):

you have to solve that equation by using "seperation of variables"

OpenStudy (baru):

\[\frac{dT_b}{dt}=-k(T_b-71)\]

OpenStudy (baru):

\[\frac{dT_b}{(T_b-71)}=-kdt\]

OpenStudy (baru):

integrate both sides